The digits Of a three digit number form a G.P.If we subtract 792 from this number whose digits are written in reverse order .If we increase the middle digit of a given number by 2, then the digits form an A.P.find numbers? 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

DEar love

let the number are ar2,ar,a

 given   ar2,ar+2 ,a  are in AP

        so 2(ar+2) = ar2 +a

             a(r-1)2 =4 .............1

and given   ar2*100 + ar*10 +a -792 = a*100 + ar*10 + ar2

                      a(r2-1)=8 .......2

from equation 1 and 2nd

 a=1 ,r=3

so number are   = 931

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin